# In-between Motion Generation Based Multi-Style Quadruped Robot Locomotion

### v1

@@ Abstract
Quadruped robots face persistent challenges in achieving versatile locomotion due to limitations in reference motion data diversity. To address these challenges, this approach introduces an in-between motion generation based multi-style quadruped robot locomotion framework, integrating synergistic advances in motion generation and imitation learning. Our approach establishes a unified pipeline addressing two fundamental aspects: First, we propose a CVAE based motion generator, synthesizing multi-style dynamically feasible locomotion sequences between arbitrary start and end states. By embedding physical constraints and leveraging joint poses based phase manifold continuity, this component produces physically plausible motions spanning multiple gait modalities while ensuring kinematic compatibility with robotic morphologies. Second, we adopt the adversarial motion priors algorithm. We validate the effectiveness of generated motion data in enhancing controller stability and improving velocity tracking performance. The proposed framework demonstrates significant improvements in velocity tracking and deployment stability. We successfully deploy the framework on a real-world quadruped robot, and the experimental validation confirms the framework’s capability to generate and execute complex motion profiles, including gallop, tripod, trotting and pacing.
@@ I INTRODUCTION
To address these issues, researchers have started to explore different approaches that can learn the style of a real dog’s movements from motion capture data. [5] encoded motion styles via latent representations, enabling bipedal robots to learn diverse movement patterns. [6, 7] utilized Adversarial Motion Priors (AMP) for imitation learning, yet remains constrained by dataset scarcity. Although these methodologies represent fundamental advances in the treatment of these issues, the core challenge of the lack of data in motion capture remains unexplored.
To overcome the data scarcity challenge, we develop an in-between motion generation algorithm specifically designed for quadruped robotic configurations for imitation learning. This framework utilizes observable data from quadruped simulation environments as both input and output, strictly adheres to joint constraints defined in the robot’s URDF model, and optimizes components of conventional motion generation methods that conflict with real-world deployment requirements. A generated motion based AMP is used to validate the effectiveness of the generated data.
Overall, these contributions form an in-between motion generation based multi-style quadruped robot locomotion framework that resolves the critical problems of reliable motion synthesis and enhanced locomotion performance. Results has been validated in the experiment.
@@ III-A1 Data Formatting
In this approach, to ensure that the generated motion adheres to the physical constraints on a real robot, we implement a differentiated input representation strategy. Specifically, the robot’s root joint is processed separately from its other joints during motion generation input construction. For the root component, we incorporate its global position within the world coordinate system (represented as 3D coordinates $[x,y,z]$) along with its rotational orientation relative to this global frame (expressed as a quaternion $[q_{x},q_{y},q_{z},q_{w}]$). Conversely, for all non-root joints, we employ joint pose (twelve joint angles) prediction rather than utilizing global positions and velocities of all physical joint points. This approach provides stronger alignment with fundamental robotic control requirements, where root state awareness and joint-space commands constitute essential control inputs. The state $s$ for one robot is shown below:
@@ III-A2 Motion Generation Network
The motion generation module of our approach is inspired by the foundational architecture proposed by [14]. While retaining the core network structure of the original model, which demonstrated efficacy in generating humanoid motion sequences, we significantly adapt the methodology to address the distinct challenges of quadruped robotic locomotion. Key modifications include the redesign of the loss functions to align with quadruped-specific implementation objectives such as joint limits and constraints, substantial alterations to the input/output representations to accommodate the quadruped robot’s observations, and the implementation of mechanisms for robust first-frame prediction crucial for deployment on physical platforms. This section details these adapted components within the inherited framework.
PAE Network There are three network needed to train in the whole training process of the motion generator. First, a periodic autoencoder (PAE) [15] is trained to generate a multi-dimensional phase manifold. The PAE network decompose motion into periodic components, obtaining an encoding phase manifold, which is a crucial element for training our motion generator.
@@
CMoEs Decoder Second, a CMoEs Decoder Network is trained to learn different local style motions, which enables the network to generate specified style motions for the motion generator. The network structure is shown in Fig. 1.
@@
$\begin{split}\mathcal{L}_{\text{KL}}=\frac{1}{2}\sum_{i=1}^{d}\left(\sigma_{i}^{2}+\mu_{i}^{2}-1-\ln(\sigma_{i}^{2})\right)\end{split}$ (2)
@@
As for the position loss, we have pos loss ($\mathcal{L}_{\text{pos}}$), joint rotation loss ($\mathcal{L}_{\text{rot}}$), orientation loss ($\mathcal{L}_{\text{ori}}$) and root position loss ($\mathcal{L}_{\text{root}}$):
@@
where $\mathbf{P}$ is the 3D global joint point calculated by forward kinematic from $s$, $\bm{\theta}$ is the rotations of robot joints, $\mathbf{R}$ is the root rotation under the global coordinate (6D representation), and $\mathbf{r}$ is the root position. $N$ is the Batch size. This loss design enables dual-constraint optimization: enforcing global coordinate requirements while maintaining robot-specific control requirements. Crucially, it incorporates penalties $\mathcal{L}_{\text{joint limit}}$ for joint limit violations when predicted motions exceed hardware capabilities. Finally we have:
@@ III-B Adversarial Imitation Learning based robot Control
Action Generation and Motion Initialization Mechanism
The policy outputs joint position deltas, and the executed joint pose is obtained by superimposing the policy-generated delta onto a predefined default configuration:
@@
Reward Function
Our reward function comprises three components: task reward $r_{t}^{T}$, regularization reward $r_{t}^{R}$, and style reward $r_{t}^{S}$, and the complete reward formulation is summarized in Table II.
@@
$||\bm{\tau}||^{2}$
$||\bm{a}_{t}-\bm{a}_{t-1}||^{2}$
$||\bm{\ddot{q}}||^{2}$
$||\max\left[0,\left|\bm{\tau}\right|-\bm{\tau^{lim}}\right]||^{2}$
@@ IV-B Imitation Learning Results Analysis
(a) Gallop
(b) Tripod
Figure 4: Real-World Deployment Result
@@ V CONCLUSIONS
In summary, we propose a in-between motion generation based multi-style quadruped robot locomotion framework consisting of two core component: a motion generator and an imitation learning network. The motion generator synthesizes reliable motion data and implements it through imitation learning methodologies. Experimental results demonstrate that this approach significantly enhances quadrupedal locomotion performance across velocity tracking and dynamic stability metrics, which successfully solves the lack of data problem in imitation learning.

### v2

@@ Abstract
Quadruped robots face persistent challenges in achieving versatile locomotion due to limitations in reference motion data diversity. To address these challenges, we introduce an in-between motion generation based multi-style quadruped robot locomotion framework. We propose a CVAE based motion generator, synthesizing multi-style dynamically feasible locomotion sequences between arbitrary start and end states. By embedding physical constraints and leveraging joint poses based phase manifold continuity, this component produces physically plausible motions spanning multiple gait modalities while ensuring kinematic compatibility with robotic morphologies. We train the imitation policy based on generated data, which validates the effectiveness of generated motion data in enhancing controller stability and improving velocity tracking performance. The proposed framework demonstrates significant improvements in velocity tracking and deployment stability. We successfully deploy the framework on a real-world quadruped robot, and the experimental validation confirms the framework’s capability to generate and execute complex motion profiles, including gallop, tripod, trotting and pacing.
@@ I INTRODUCTION
To address these issues, researchers have started to explore different approaches that can learn the style of a real dog’s movements from motion capture data. [5] encoded motion
(a) Gallop
(b) Tripod
Figure 1: Deployment result. Gallop and Tripod motion that learned from the in-between motion generated motion.
styles via latent representations, enabling bipedal robots to learn diverse movement patterns. [6, 7] utilized adversarial motion priors (AMP) for imitation learning, yet remains constrained by dataset scarcity. Although these methodologies represent fundamental advances in the treatment of these issues, the core challenge of the lack of data in motion capture remains unexplored.
To overcome the data scarcity challenge, we develop an in-between motion generation algorithm specifically designed for quadruped robotic configurations for imitation learning. We propose an in-between motion generation based multi-style quadruped robot locomotion framework, which utilizes observable data from quadruped simulation environments as both input and output, strictly adhering to joint constraints from robot physical capabilities. Furthermore, we redesign the loss functions, making them adapted to joint limits and constraints for implementation objectives. Moreover, a first-frame prediction is proposed for deployment. A generated motion based AMP is used to validate the effectiveness of the generated data. Results has been validated in the real-world deployment shown in Fig. 1.
@@ III-A1 Data Formatting
In this approach, to ensure that the generated motion adheres to the physical constraints on a real robot, we implement a differentiated data formatting strategy. Considering robot root stability, physical constraints, and quadruped robot dynamics, the robot’s root joint is processed separately from its other joints during motion generation input construction. We incorporate its global position within the world coordinate system $\mathbf{p}_{root}$ along with its rotational orientation relative to this global frame $\mathbf{R}_{\text{root}}$. And we employ joint pose of the quadruped robot $\mathbf{q}$ rather than utilizing global positions and velocities of all physical joint points. This approach provides stronger alignment with fundamental robotic control requirements, where root state awareness and joint-space commands constitute essential control inputs. The state $s$ for one robot is shown below:
@@ III-A2 Motion Generation Network
The motion generation module of our approach is inspired by the foundational architecture proposed by [14]. While retaining the core network structure of the original model, which demonstrated efficacy in generating humanoid motion sequences, we significantly adapt the methodology to address the distinct challenges of quadruped robotic locomotion. Our motion genetator composed three net: PAE network, CMoEs network, and sampler network.
PAE Network First, a periodic autoencoder (PAE) [15] is trained to generate a multi-dimensional phase manifold. The PAE network decompose motion into periodic components, obtaining an encoding phase manifold, which is a crucial element for training our motion generator.
@@
CMoEs Decoder CMoEs Decoder is used to learn different local style motions, which enables the network to generate specified style motions for the motion generator. The network structure is shown in Fig. 2.
@@
$\mathcal{L}_{\text{KL}}=\frac{1}{2}\sum_{i=1}^{d}\left(\sigma_{i}^{2}+\mu_{i}^{2}-1-\ln(\sigma_{i}^{2})\right)$ (2)
@@
As for the position loss, we have pos loss ($\mathcal{L}_{\text{pos}}$), joint rotation loss ($\mathcal{L}_{\text{rot}}$), orientation loss ($\mathcal{L}_{\text{ori}}$) and root position loss ($\mathcal{L}_{\text{root}}$), which is shown in Eq. 4, where $\mathbf{P}$ is the 3D global joint point calculated by forward kinematic from $s$, $\bm{\theta}$ is the rotations of robot joints, $\mathbf{R}$ is the root rotation under the global coordinate (6D representation), $\mathbf{r}$ is the root position, and $N$ is the Batch size. This loss design enables dual-constraint optimization: enforcing global coordinate requirements while maintaining robot-specific control requirements. Crucially, it incorporates penalties $\mathcal{L}_{\text{joint limit}}$ for joint limit violations when predicted motions exceed hardware capabilities. Finally we have:
@@ III-B Adversarial Imitation Learning based robot Control
Action Generation and Motion Initialization Mechanism The policy outputs joint position deltas, and the executed joint pose is obtained by superimposing the policy-generated delta onto a predefined default configuration:
@@
Reward Function Our reward function comprises three components: task reward $r_{t}^{T}$, regularization reward $r_{t}^{R}$, and style reward $r_{t}^{S}$, and the complete reward formulation is summarized in Table II.
@@
$\bm{\tau}||^{2}$
$\bm{a}_{t}-\bm{a}_{t-1}||^{2}$
$\bm{\ddot{q}}||^{2}$
$\max\left[0,\left|\bm{\tau}\right|-\bm{\tau^{lim}}\right]||^{2}$
@@ V CONCLUSIONS
In summary, we propose an in-between motion generation based multi-style quadruped robot locomotion framework, which is capable of generating motions at arbitrary velocities using sparse motion capture data, while achieving accurate velocity tracking performance through imitation learning based deployment on real-world robots. Experimental results demonstrate that this approach significantly enhances quadrupedal locomotion performance across velocity tracking and dynamic stability metrics, which successfully solves the lack of data problem in imitation learning.